Liquid hexane will react with gaseous oxygen to produce gaseous carbon dioxide and gaseous water . Suppose 4.3 g of hexane is mixed with 23.0 g of oxygen. Calculate the maximum mass of carbon dioxide that could be produced by the chemical reaction. Be sure your answer has the correct number of significant digits.
1
Expert's answer
2020-12-06T13:51:38-0500

2C6H14 + 19O2 = 12CO2 + 14H2O

M (CO2) = 44.01 g/mol

M (O2) = 32 g/mol

M (C6H14) = 86.18 g/mol

n (C6H14) = 4.3 /86.18 = 0.05 mol

n (O2) = 23.0 /32 = 0.72 mol

Which reactant is limiting? C6H14 = 0.05/2 = 0.025; O2 = 0.27/19 = 0.04

Thus C6H14 is a limiting reactant.

n (CO2) that can be produced = 0.05/2 x 12 = 0.3 mol

m (CO2) = 0.3 mol x 44.01 g/mol = 13.2 g (to 3 sig. figs.)